### v1

# Steer-Bench: A Benchmark for Evaluating the Steerability of Large Language Models

Abstract
Steerability, or the ability of large language models (LLMs) to adapt outputs to align with diverse community-specific norms, perspectives, and communication styles, is critical for real-world applications but remains under-evaluated. We introduce Steer-Bench, a benchmark for assessing population-specific steering using contrasting Reddit communities. Covering 30 contrasting subreddit pairs across 19 domains, Steer-Bench includes over 10,000 instruction-response pairs and validated 5,500 multiple-choice question with corresponding silver labels to test alignment with diverse community norms. Our evaluation of 13 popular LLMs using Steer-Bench reveals that while human experts achieve an accuracy of 81% with silver labels, the best-performing models reach only around 65% accuracy depending on the domain and configuration. Some models lag behind human-level alignment by over 15 percentage points, highlighting significant gaps in community-sensitive steerability. Steer-Bench is a benchmark to systematically assess how effectively LLMs understand community-specific instructions, their resilience to adversarial steering attempts, and their ability to accurately represent diverse cultural and ideological perspectives.111Code and data are available at https://github.com/kaichen23/steer-bench.
Steer-Bench: A Benchmark for Evaluating the
Steerability of Large Language Models
Kai Chen1,2, Zihao He1,2, Taiwei Shi1, Kristina Lerman2 1Department of Computer Science, University of Southern California 2Information Sciences Institute, University of Southern California {kchen035, zihaoh, taiweish}@usc.edu, lerman@isi.edu
1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in following instructions, generating coherent text, and adapting to various contexts Lou et al. (2024); Tam et al. (2024); Zhang et al. (2023). A key dimension of these capabilities is steerability—the ability of an LLM to tailor its outputs in response to specific guidance, constraints, or norms provided by users or developers.
Among these capabilities, steerability has emerged as an important dimension for practical applications. Steerability refers to a large language model’s ability to adapt its outputs according to specific guidance, preferences, norms, or constraints provided by users or developers. Figure 1 illustrates steering LLMs using instruction-response pairs. Steerability encompasses multiple related behaviors. At its most basic, it includes instruction following: generating outputs that conform to explicit prompts. More advanced forms include model steering, where outputs are conditioned on explicit attributes or requirements; role-playing, in which models adopt consistent personas or communication styles Shanahan et al. (2023); Wang et al. (2024a); and personalization, where outputs reflect individual user preferences Li et al. (2024b); Kumar et al. (2025). Another essential behavior for many real-world applications is norm following—i.e., adhering to the values, rules, and discourse patterns of particular communities Shi et al. (2024b); Liu et al. (2025); Li et al. (2024a).
While numerous benchmarks exist to evaluate general LLM capabilities, such as reasoning, reading comprehension Rajpurkar et al. (2016); Dua et al. (2019), instruction following Ouyang et al. (2022), ethical alignment Hendrycks et al. , and personalization Kumar et al. (2025), few are explicitly designed to evaluate higher-order capabilities like steerability. This limits our ability to systematically measure model alignment with diverse social, cultural, or ideological contexts.
To address this gap, we introduce Steer-Bench, a benchmark designed to evaluate population-specific steering in LLMs. Steer-Bench draws on Reddit, a platform comprising thousands of topic-specific communities, many of which represent contrasting perspectives on shared issues (e.g., r/liberals vs. r/conservatives, r/parenting vs. r/Childfree, r/Linux vs. Windows). These forums provide natural discourse that reflects distinct norms, rhetorical styles, and worldviews of distinct communities, making Reddit an ideal setting for assessing whether models can produce outputs that are not only coherent but also contextually appropriate across different social or cultural contexts.
Steer-Bench includes 30 contrasting subreddit pairs across 19 domains, from which we curated over 10,000 open-ended question-answer pairs that reflect diverse community viewpoints. The benchmark also features a set of multiple-choice questions with corresponding silver labels, validated by human annotators, to provide high-quality ground truth for alignment evaluation.
We use Steer-Bench to evaluate the steerability of 13 popular open-source and proprietary LLMs. We show that steerability improves with model size in both in-context learning and instruction-tuning settings. Furthermore, we find that contextualized instructions significantly enhance steering performance relative to relying solely on pretrained knowledge. Finally, we demonstrate that model families exhibit distinct sensitivities to steering across domains, underscoring the importance of training methodology and architecture in enabling subpopulation alignment.
Our key contributions are:
We introduce STEER-BENCH, a benchmark for evaluating community-specific steerability in LLMs across 30 subreddit pairs and 19 domains, covering over 10,000 instruction-response pairs and 5,500 multiple-choice questions grounded in contrasting online communities.
We evaluate 13 popular LLMs under both in-context learning and supervised finetuning settings, revealing that steerability improves with model scale and contextualized prompting, but varies significantly by model family and domain.
We identify substantial performance gaps between models and human-aligned labels—especially in ideologically sensitive domains—highlighting the challenges of aligning LLMs with diverse cultural and social norms.
2 Related Work
Instruction following
Instruction following has become foundational for large language models, encompassing techniques like instruction tuning Ouyang et al. (2022); Wang et al. (2023) and in-context learning Zhao et al. (2025); Edwards and Camacho-Collados (2024). Recent research has proposed comprehensive benchmarks and frameworks to systematically evaluate the proficiency of LLMs in executing instructed tasks Zeng et al. (2024); Zhou et al. (2023). These benchmarks critically assess approaches aimed at improving LLMs’ ability to interpret and adhere to complex instructions involving multiple constraints He et al. (2024a); Sun et al. (2024); Zhang et al. (2024). Additionally, Tam et al. (2024) investigates the impact of structured output constraints on LLM performance, providing further insights into factors influencing instruction-following effectiveness.
Model steering
Researchers have explored various methods for steering the behavior of large language models, seeking to enhance utility, controllability Bayat et al. (2025); Guo et al. (2024), and alignment with human preferences Alves et al. (2023). Dong et al. (2023) introduce SteerLM, which conditions responses on explicit attributes during inference, providing a supervised fine-tuning alternative to RLHF for direct behavioral control. Similarly, Cao et al. (2024) creates effective steering vectors through bi-directional preference optimization that precisely influence generation probabilities based on human preference pairs. However, challenges remain: Chen et al. (2024a) reveals LLMs’ susceptibility to ideological steering through minimal biased data exposure, Santurkar et al. (2023) finds that attempts to steer LLMs toward specific demographic viewpoints through prompting resulted in only modest improvements, indicating difficulties in effectively aligning models with diverse human opinions. Our work focuses on systematically measuring the steering capabilities of large language models.
Role-play
The role-playing capabilities of LLMs have emerged as a popular focus of research, resulting in the development of diverse frameworks and systematic methodologies for enhancing and evaluating these capabilities. Shanahan et al. (2023) conceptualizes LLMs as role-players using folk psychological terms without anthropomorphizing them, while Chen et al. (2024b) provides a comprehensive taxonomy covering data, models, alignment, and evaluation challenges. He et al. (2024c) and Santurkar et al. (2023) measure the opinion and affective alignment of LLMs to different social groups when being steered to mimic them. Several frameworks have been proposed to enhance role-playing capabilities. Wang et al. (2024a) introduce RoleLLM with the RoleBench dataset; Lu et al. (2024) develop Ditto for maintaining consistent role identities through dialogue simulation; Ran et al. (2024) create ROLEPERSONALITY to incorporate psychological dimensions into character modeling. Similarly, Shao et al. (2023) propose Character-LLM, which simulates specific individuals by incorporating their profiles and emotional states.
Synthetic Data
The scarcity of high-quality, diverse, and labeled datasets has long presented a bottleneck for training effective Large Language Models. To address this, researchers have increasingly turned to synthetic data generation as a viable alternative or supplement to real-world data Wang et al. (2024b); Zhong et al. (2025); Ghanadian et al. (2024); Hämäläinen et al. (2023). For example, Wang et al. (2023); Xu et al. (2024); Li et al. (2024c) generate complex instruction data from simple seed instructions using large language models. Similarly, He et al. (2024b); Shi et al. (2024a) generate instruction data based on massive noisy social media data, while Chen et al. (2024a) produces high-quality synthetic instruction-response pairs from political surveys.
3 Problem Definition
A topical domain (e.g., gender, technology, or religion) comprises contrasting communities , that express divergent perspectives and rhetorical styles. Each community generates a text corpus (e.g., Reddit posts and comments) that captures its mindset, communication norms, and ideological stance. Our objective is to evaluate the ability of a large language model to be steered towards the perspective of each community , such that its outputs reflect the community’s distinctive voice and beliefs.
To steer a model towards a target community , we provide it with a set of demonstrations (instruction-response pairs) that exemplify the community’s responses to open-ended instructions. These demonstrations can be used either as in-context examples or as finetuning data. When guided by , the steered model is expected to produce outputs that align with ’s ideology and communication style222In-context learning does not modify the model weights, thus not leading to a different model. However, we still denote the model as to signify the effect of steering..
To systematically assess steerability across a wide range of community perspectives, we construct Steer-Bench, a benchmark of 30 contrasting subreddit pairs across 19 domains. For each pair, we automatically generate both instruction-response demonstrations and multiple-choice questions using an advanced LLM (GPT-4o), based on topic-specific samples from the community corpora and . The multiple-choice questions serve as structured evaluations that measure how well a model steered toward community or selects the answer consistent with that community’s perspective. This setup allows us to benchmark and compare the steerability of different LLMs under both in-context and finetuning-based steering paradigms.
4 Steer-Bench Construction
To evaluate the steerability of LLMs toward specific community perspectives, we construct Steer-Bench, a benchmark comprising automatically generated steering demonstrations and evaluation instances derived from contrasting online communities. The construction pipeline of Steer-Bench is shown in Figure 2. In this section, we describe how we identify community pairs, collect data, generate instruction-response demonstrations for steering models, and build multiple-choice evaluation instances for assessing whether a model steered toward a community accurately reflects its views.
4.1 Community and Domain Selection
Reddit is a social media platform composed of numerous communities, known as subreddits, each dedicated to discussions around specific topics or themes. These communities facilitate interactions through user-generated posts and comments, providing a rich environment for members to express diverse opinions and ideologies Chen et al. (2023).
We define a topical domain (e.g., politics or diet) as a set of contrasting communities , that engage in discussion around a shared theme but from ideologically distinct perspectives. Each community produces a domain-specific corpus through user-generated Reddit submissions and comments. We curate 30 such community pairs across 19 domains, including gender (e.g., r/AskWomen vs. r/AskMen), religion (e.g., r/atheism vs. r/Christianity), and technology (e.g., r/apple vs. r/Android), etc. These contrasting pairs are selected based on domain expertise and LLM-assisted analysis.
We collect Reddit data, including both submissions and comments, for each community over the course of 2024 using Academic Torrent.333https://academictorrents.com/ The recency of data helps minimize overlap with the pretraining data of the current large language models. Submissions and comments are treated uniformly as individual documents. We filter out submissions with fewer than 5 comments, remove moderated comments, and treat each submission or comment as a document in . To reduce bias from imbalanced community sizes, we subsample up to 500,000 documents per community and cap sampling from the larger community in each pair at 3 the size of the smaller one. Full statistics for all subreddit pairs and domains are shown in Table 1 in Appendix B.
4.2 Topic Identification
To generate meaningful and comparable demonstrations across contrasting communities, we first identify overlapping topics within each subreddit pair. We concatenate and apply BERTopic Grootendorst (2022) to extract a list of topics . For each topic , we retain it if both communities contribute at least 200 comments, ensuring that perspectives from both and are represented. We select up to the top 20 such topics per pair where each topic is discussed in both subreddits, yielding a diverse and balanced set of discussion themes for data generation. The identified topics is shown in Table 2 to 9 in Appendix B.
4.3 Instruction-Response Generation
Community-Cross-Instruct He et al. (2024b) automatically generates instruction-response pairs and multiple question-answer pairs from social media data; COMPO Kumar et al. (2025) leverages contrasting subreddit norms for personalized preference optimization in language models. Built upon these two frameworks, we introduce a methodology that automatically generates benchmark datasets from social media data to systematically evaluate the steerability of LLMs.
We generate a set of community-aligned instruction-response pairs for each using GPT-4o as a synthetic data generator. For each topic , we sample 50 comments from and , anonymize subreddit names as “r/A” and “r/B”, and prompt GPT-4o to generate three instructions {, that elicit contrasting viewpoints across the two communities. See Figure 5 in Appendix A for the prompting template. This anonymization encourages GPT-4o to generate instruction-response pairs based strictly on the provided comments rather than the prior knowledge of community. For each instruction GPT-4o generates two responses and aligned with and , respectively. This yields paired demonstrations , suitable for either supervised finetuning or in-context learning.
To ensure diversity, we repeat this procedure four times per topic with different document samples, thus leading to open-ended instruction-response pairs per topic. In addition, to better profile a community, we generate single-community instruction-response pairs for each by targeting community-specific topics (ones that are not shared with the other community in the pair). The prompting template is shown in Figure 6 in Appendix A. These examples are used exclusively for supervised finetuning (5.5).
4.4 Question-Answer Generation
To measure steerability in a structured and automated manner, we generate multiple-choice questions , where each question probes a topic and is the answer choice aligned with either or . GPT-4o is prompted to create such a question and select the correct answer and , aligned with and respectively. The prompting template is shown in Figure 5 Appendix A. In practice, GPT-4o generates two questions { per iteration, along with the three instructions {, for better efficiency. For each topic , we repeat the generation four times.
These questions act as synthetic surveys to test whether a model steered toward can produce responses aligned with . We refer to the GPT-4o answer as the silver label. Questions with ambiguous or invalid answers, such as “not discussed”, “no specific mention”, and “unclear”, are discarded.
4.5 Dataset Statistics
The final Steer-Bench benchmark contains: (1) 8,328 instruction-response pairs derived from 30 subreddit pairs and 347 shared topics; (2) 5,552 multiple-choice questions with corresponding silver labels; (3) 3,285 additional single-community instructions solely for supervised finetuning.
Each paired instruction contributes two demonstrations—one per community—and similarly, each multiple-choice question contributes two question-answer pairs. Domain coverage is visualized in Figure 12 in Appendix B and full statistics are listed in Table 10 in Appendix B.
We observe that r/Gender, r/Politics and r/Gaming have the largest representation, which feature prominently in online discourse and contain well-established contrasting communities with distinct perspectives. The variation in topic count of different subreddit pairs reflects the breadth of discussion and the availability of comparable content across contrasting subreddits.
4.6 Human Validation
To validate that generated demonstrations and questions faithfully represent community views, we conduct a human annotation study using a pool of four annotators familiar with Reddit culture. Annotators assign community labels to model-generated responses and select the correct answers to multiple-choice questions. The annotators volunteered for this task with full awareness that their annotations would only be used to evaluate the performance of GPT-4o’s generation.
The evaluation is conducted via Google Forms, with 30 sections per form, each for a randomly sampled topic from a subreddit pair. For each section, we sample two instructions and one multiple-choice question from the same topic. Each section includes four questions: (1) whether the annotator is familiar with the domain and two subreddits, (2)-(3) which response corresponds to a specific subreddit for each instruction, and (4) a multiple-choice question where annotators select the correct answer from two options. An example section in the form is shown in Figure 13 in Appendix C.
After collecting responses from all four annotators, we filter out sections where annotators indicated unfamiliarity with the domain or subreddits. The agreement between four human annotators is 0.712 measured by Fleiss’ Kappa. Golden labels are generated by soft voting from four human annotators with annotator confidence scores (score 1 for "Yes" and 0.5 for "Maybe" under the first question in the section). The inter-rater agreement between the golden labels and silver labels (GPT-4o generated answers) is 0.815 measured by Cohen’s Kappa.
5 Evaluating LLM Steerability
5.1 Experiments
We evaluate the steerability of LLMs using the Steer-Bench benchmark. Given a target community with associated corpus , our goal is to determine how effectively a language model can be steered towards such that its outputs align with the community’s perspectives and beliefs. Steering is accomplished by conditioning the model on a set of demonstrations , where each is an open-ended instruction and is a response reflective of ’ ideology.
Although subreddit pairs (, ) are used during data construction to identify shared topics across communities, each community is treated independently during evaluation. That is, for each community , we steer the model using instruction-response pairs and evaluate it solely on question-answer pairs relevant to that community.
We evaluate steerability under two settings: (1) in-context learning, where demonstrations are provided as examples in the prompt at inference time; and (2) supervised finetuning, where the model is explicitly finetuned on the demonstrations.
5.2 Experimental Setup
Our experiments span 13 LLMs, covering both open-source and proprietary families:
Open-source: Llama-3 (3B, 8B, 70B), Qwen-2.5 (3B, 7B, 14B, 32B, 72B), Mistral-7B, DeepSeek-v3
Proprietary: Claude-3.5-Haiku, Claude-3.7-Sonnet, GPT-4o-Mini
All models are instruction-tuned variants intended for conversational interaction and controllable generation. Details for each model are provided in Table 13 in Appendix D.
For in-context learning, for a multi-choice question on topic , we retrieve the 12 on-topic demonstrations , and embed them as in-context examples in the prompt. We use vLLM to query each model, with a temperature of 0.75 and top-p of 0.9. For supervised finetuning, for each community, we finetune a representative subset of open-weight models: Llama-3 (3B, 8B) and Qwen-2.5 (3B, 7B, 14B), on . We perform full-parameter training for 2 epochs with a batch size of 8. The learning rate is set to 8e-6 for 3B models and 6e-6 for larger ones. Training is conducted on 8 NVIDIA H100 GPUs.
5.3 Evaluation Protocol
Given a steered (via in-context or finetuning), we present it with a set of community-specific multi-choice questions . Each question targets a topic discussed by community , and is the answer (silver label) generated by GPT-4o based on . We measure the accuracy as the proportion of model responses that match the silver labels. This quantifies how well reflects community-aligned perspectives when steered towards .
5.4 Steering via In-context Learning
We assess five prompting configurations to evaluate the effectiveness of in-context learning:
1.
Vanilla: Answer multi-choice question without context;
- 2.
- 3.
- 4.
- 5.
Figure 3 presents results of the steerability evaluation of different models across five configurations. For nearly all models, steering using in-topic few-shot demonstrations (Config 4) outperforms those under baseline conditions (Configs 1 and 2). This demonstrates that explicitly contextualizing a model with the community’s own responses is more effective than relying on prior knowledge alone. Further improvements are observed under Config 5, particularly for Qwen and Llama, where the subreddit identifier complements in-topic examples by providing a stronger community grounding.
Claude-3.5-Haiku and Sonnet deviate from this trend, achieving their highest performance in Config 3 (Subreddit Identifier). This suggests Claude’s pretraining may already encode subreddit-specific priors more effectively than other models. Mistral-7B, by contrast, shows large gains with in-topic demonstrations (Config 4), highlighting its dependence on prompt conditioning over internal knowledge. Notably, DeepSeek-v3 fails to benefit from any configuration, achieving uniformly low performance, suggesting poor alignment with English-speaking community norms—possibly due to its predominantly non-English pretraining.
Across families, we observe strong within-family scaling. For instance, steerability improves steadily from Qwen-2.5-3B to 72B and from Llama-3.2-3B to 70B, confirming that larger models are better able to absorb and adapt to contextual cues. However, cross-family performance gaps persist. For example, Claude-3.5-Haiku and GPT-4o-mini outperform several larger open-weight models, indicating that architectural and pretraining design are at least as important as scale.
5.4.1 Domain-level Steerability
Figure 4 presents model accuracy by domain using in-context learning (“In-topic Few-shot”, Config 4), highlighting cross-domain variation in steerability. More detailed performance is shown in Tables 14 and 15 in Appendix. These results reveal important differences in how LLMs engage with different domains.
Easy domains
Diet, Self-Improvement, and Religion emerge as domains where nearly all models perform well, with top models achieving >0.70 accuracy. These domains feature clearly articulated community norms (e.g., keto vs. vegan, GetMotivated vs. getdisciplined) and strongly polarized rhetoric, which likely facilitates easier identification of community-aligned answers.
Challenging domains
Music and Technology exhibit lower performance across all models. These domains may be harder because they feature preferences rather than ideologies, with fewer stylistic or conceptual anchors to distinguish perspectives (e.g., electronic vs. classical music).
Ideologically Sensitive Domains
Domains like Abortion, Politics, and Social Issues show large inter-model disparities. Qwen-2.5-32B achieves 0.812 in Abortion, while Llama-3.2-3B scores only 0.281—an enormous 53-point gap. This suggests that controversial domains demand greater contextual understanding and capacity to model ideological positions—something only a subset of models currently achieve well.
Fine-Grained Specialization
Model families exhibit distinct areas of strength: Qwen excels in Abortion and Finance, while Llama-3.3-70B dominates Education and Science. Claude performs well in Technology and News but underperforms in Diet.
Scaling Trends
Within nearly every domain, model performance increases with size. For instance, Qwen models improve monotonically across most domains from 3B to 32B before plateauing or slightly declining at 72B, possibly due to overfitting or prompt length limitations. Llama-3.2-3B performs poorly across the board, but its 70B counterpart ranks among the top in 10+ domains, showing that capacity is necessary for nuanced community modeling.
5.5 Steering via Supervised Finetuning
We evaluate model steerability via supervised finetuning. Overall, in-context learning outperforms finetuning across most models, suggesting that prompt-based steering is more effective given our current data scale. However, smaller models (e.g., Llama-3.2-3B) benefit more from finetuning, and steerability still improves with model size. Detailed comparisons and analysis are provided in Appendix E.
6 Conclusion
STEER-BENCH provides a framework for evaluating how effectively LLMs adapt to diverse community perspectives. Our assessment of 13 models shows a significant gap between human performance and even the best LLMs. Steerability improves with model size and contextual demonstrations but varies considerably across domains, with models performing better in communities with clear norms than in subjective domains.
These findings are critical as LLMs become embedded in real-world applications where cultural sensitivity matters. STEER-BENCH enables more precise evaluation of community alignment and supports development of models that better respect diverse social contexts. Future work should expand this approach to multimodal content, multilingual communities, and dynamic feedback systems.
Limitations
Reddit-centric community coverage
Steer-Bench is built entirely on Reddit data, which limits its scope to communities active on that platform. Reddit users tend to be English-speaking, relatively tech-savvy, and concentrated in certain regions and age groups. As a result, the benchmark may not generalize to populations that are less active online or are better represented on other platforms, such as X, Weibo, TikTok, or regional forums.
Bias in GPT-4o-generated supervision
Instruction-response pairs and multiple-choice questions are generated using GPT-4o, which may introduce biases stemming from its own pretraining data and alignment procedures. Although validated by human annotators, the silver labels used for evaluation reflect GPT-4o’s interpretations of community views, which may not always faithfully capture the true diversity or nuance within each community. This also risks favoring models that resemble GPT-4o over those trained differently.
Binary community framing
Each domain in Steer-Bench is represented as a binary contrast between two communities (e.g., r/Parenting vs. r/Childfree). While this helps isolate divergent viewpoints, it oversimplifies many ideological or cultural landscapes, which often span a continuum of perspectives. The current structure may miss important intra-community variation or overlook more subtle ideological gradients.
Simplified evaluation format
Evaluation relies on multiple-choice questions with a single correct answer per question. This structure facilitates automated scoring but does not account for the complexity, ambiguity, or subjectivity inherent in community-specific responses. Some questions may have more than one plausible answer depending on interpretation, and models may produce valid but non-matching responses that are penalized under this scheme.
Ethics Statement
Use of publicly available data
This work relies exclusively on publicly available Reddit data collected through Academic Torrent. Reddit users post content under pseudonyms, and our data collection excludes any personally identifiable information. In addition, we do not attempt to deanonymize users or link posts across communities. All preprocessing and filtering procedures are designed to protect user privacy and aggregate content at the community level.
Respect for community norms
Although Reddit content is public, many communities have distinct values, expectations, and sensitivities. When sampling and analyzing posts, we took care to anonymize subreddit names during prompt generation and to avoid making judgments about the correctness or desirability of any community’s views. Our benchmark is intended to evaluate LLMs’ ability to reflect community perspectives, not to endorse or critique them.
Bias and harm considerations
Some of the subreddit pairs in our benchmark engage with politically charged, ideologically sensitive, or potentially harmful topics (e.g., abortion, religion, gender). While our methodology is designed to surface differences in rhetorical style and viewpoint, models may inadvertently learn, reinforce, or amplify biases present in the data. We urge caution in downstream use and recommend further auditing before deployment in high-stakes settings.
Intended use and limitations
Steer-Bench is intended for research purposes only. It should not be used to develop or deploy systems that impersonate real individuals, simulate communities without transparency, or manipulate public opinion. The benchmark is a tool for studying steerability in controlled conditions—not for operationalizing sensitive sociocultural behaviors in production systems.
Acknowledgments
This project was funded in part by DARPA under contract HR001121C0168.
Appendix A Prompting Templates
Appendix B Data statistics
Subreddit_pair | No. comments |
Liberal_Conservative | 143,277 |
democrats_republicans | 42,692 |
AskALiberal_AskConservatives | 483,627 |
news_conspiracy | 490,366 |
atheism_Christianity | 499,304 |
exmuslim_islam | 476,219 |
AskWomen_AskMen | 485,759 |
Feminism_MensRights | 287,804 |
AskFeminists_MensLib | 114,466 |
abortion_prolife | 186,486 |
personalfinance_wallstreetbets | 489,798 |
apple_Android | 425,302 |
linux_windows | 200,278 |
Parenting_childfree | 489,408 |
keto_vegan | 497,936 |
carnivore_vegetarian | 44,210 |
realmadrid_Barca | 496,762 |
warriors_lakers | 499,840 |
xbox_playstation | 495,315 |
leagueoflegends_DotA2 | 491,964 |
simpleliving_UnethicalLifeProTips | 499,531 |
environment_climateskeptics | 91,210 |
electricvehicles_regularcarreviews | 499,692 |
GetMotivated_getdisciplined | 166,963 |
DecidingToBeBetter_howtonotgiveafuck | 95,001 |
Teachers_homeschool | 346,713 |
antiwork_WorkReform | 499,366 |
science_philosophy | 185,371 |
science_askphilosophy | 349,502 |
Subreddit Pair | Topic Index | Topic Keywords |
AskWomen_AskMen | 0 | money, pay, debt, jobs, college, income, career, rich, savings, retirement |
1 | she, was, relationship, with, friend, did, back, for, it, but | |
2 | age, older, women, gap, young, dating, mature, 30s, olds, attractive | |
4 | single, dating, meet, life, alone, relationships, date, yourself, be, want | |
5 | he, ex, we, relationship, with, ended, back, wasn, time, friend | |
7 | movies, watched, show, anime, characters, horror, tv, scenes, episodes, love | |
8 | attractive, ugly, looks, personality, attraction, appearance, unattractive, beauty, attractiveness, women | |
9 | apps, dating, matches, tinder, bumble, meet, profiles, date, swiping, looking | |
10 | mom, parents, sister, mother, father, family, he, mum, relationship, son | |
12 | income, pay, finances, financially, accounts, rich, women, earns, split, expenses | |
13 | cheese, pizza, eat, meat, add, cooked, salad, delicious, toast, taste | |
15 | friends, female, friendship, friend, platonic, friendships, romantic, attracted, men, relationship | |
16 | men, they, women, are, lack, man, things, woman, all, be | |
17 | compliment, compliments, complimented, complimenting, smile, shirt, say, handsome, told, men | |
18 | smoking, addiction, smoked, cigarettes, nicotine, vape, smokes, cannabis, addict, shrooms | |
19 | straight, bi, lesbian, bisexual, sexuality, lesbians, friends, queer, heterosexual, lgbt | |
Feminism_MensRights | 0 | trump, biden, left, voting, election, democrats, kamala, conservative, voters, white |
1 | feminism, feminists, feminist, equality, rights, movement, women, are, issues, them | |
2 | sub, post, mods, ban, subreddits, on, rights, about, comments, feminist | |
3 | war, military, conscription, drafted, men, combat, selective, wars, are, equality | |
4 | islam, muslim, religions, hijab, muslims, bible, is, islamic, wear, women | |
5 | abortion, abortions, fetus, rights, choice, birth, abort, states, roe, reproductive | |
6 | her, she, was, lawyer, and, with, told, didn, ex, police | |
7 | marriage, divorce, married, alimony, marry, marriages, prenup, relationship, wife, divorced | |
8 | bear, choose, question, forest, woman, encounter, men, alone, be, safer | |
9 | sex, sexual, orgasm, virgin, count, body, women, orgasms, high, having | |
11 | chores, sahm, house, husband, wife, household, housework, job, laundry, tradwife | |
13 | bisexual, lesbians, homophobia, lgbt, are, queer, sexuality, homophobic, gays, friends | |
14 | india, indian, rape, cases, indians, dowry, are, law, women, violence | |
17 | rape, sexual, victims, penetrate, rapists, survey, cdc, data, women, report | |
18 | workers, prostitution, trafficking, industry, prostitutes, is, prostitute, exploitation, job, trafficked | |
AskFeminists_MensLib | 0 | men, feminism, women, of, are, feminist, it, patriarchy, feminists, all |
1 | trump, vote, party, they, left, biden, voting, democrats, white, election | |
2 | rape, violence, abuse, victims, the, are, and, women, victim, assault | |
3 | this, it, what, re, just, is, comment, post, read, people | |
4 | boys, education, jobs, are, women, teachers, male, is, pay, schools | |
6 | sex, consent, sexual, orgasm, partner, not, women, want, if, can | |
7 | body, fat, beauty, and, people, weight, surgery, appearance, eating, standards | |
8 | military, war, draft, conscription, ukraine, women, drafted, it, are, soldiers | |
9 | characters, character, female, movie, show, it, is, strong, series, woman | |
10 | emotions, emotional, anger, emotion, feelings, men, cry, be, are, can | |
13 | masculinity, masculine, feminine, femininity, positive, traits, be, men, gender, toxic | |
17 | people, bad, it, moral, do, of, be, are, think, morality |
Subreddit Pair | Topic Index | Topic Keywords |
AskALiberal_ AskConservatives | 0 | abortion, abortions, fetus, choice, roe, medical, is, autonomy, bodily, murder |
1 | border, immigration, immigrants, asylum, illegals, undocumented, immigrant, deport, deportation, wall | |
2 | gun, guns, shootings, firearms, weapons, amendment, firearm, laws, militia, rifles | |
3 | israel, hamas, palestinians, palestinian, gaza, palestine, genocide, israelis, conflict, hostages | |
4 | religious, god, christianity, christians, bible, religions, commandments, catholic, believe, atheist | |
5 | ukraine, russia, nato, putin, war, ukrainians, russians, crimea, invaded, nuclear | |
6 | schools, education, students, teachers, degree, teacher, debt, funding, teaching, colleges | |
7 | she, vp, has, candidate, think, that, biden, president, of, would | |
8 | healthcare, insurance, medicare, aca, companies, medicaid, government, countries, us, premiums | |
9 | biden, trump, candidate, voters, party, democrats, election, joe, 2020, democratic | |
10 | harris, trump, biden, vote, candidate, voters, win, election, democrats, democratic | |
11 | comment, conversation, op, response, point, reply, understand, argument, thread, discussion | |
12 | taxes, spending, deficit, cut, capital, increase, rates, wealthy, taxing, government | |
13 | covid, vaccines, pandemic, vaccinated, flu, vaccination, deaths, lockdowns, polio, measles | |
14 | news, media, cnn, sources, msnbc, npr, propaganda, journalism, journalists, unbiased | |
15 | capitalism, socialism, communism, socialist, capitalist, marx, marxism, marxist, democracy, economy | |
16 | trans, gender, puberty, dysphoria, transgender, medical, transitioning, hormones, minors, therapy | |
17 | kamala, she, harris, trump, biden, candidate, primary, voters, election, democrats | |
18 | protests, insurrection, riots, riot, rioters, peaceful, protestors, protesters, capital, election | |
19 | mods, banned, conservatives, post, askconservatives, subreddits, rules, liberal, askaliberal, moderation | |
Liberal_ Conservative | 0 | she, was, be, it, has, like, with, if, trump, what |
1 | ukraine, russia, war, putin, israel, nato, hamas, iran, are, military | |
2 | court, he, case, is, president, supreme, judge, law, trial, immunity | |
3 | biden, trump, debate, in, president, vote, has, it, joe, who | |
4 | border, immigration, bill, immigrants, asylum, they, mexico, in, is, legal | |
5 | kamala, biden, harris, for, they, has, vote, as, debate, if | |
6 | left, party, democrats, are, right, republicans, conservatives, liberal, conservative, liberals | |
7 | abortion, abortions, birth, women, is, roe, ivf, babies, issue, pregnant | |
8 | blue, state, live, california, cities, states, rural, texas, move, liberal | |
9 | harris, biden, to, campaign, vote, that, if, they, has, win | |
11 | gun, guns, firearms, amendment, shootings, ban, assault, laws, mass, carry | |
12 | christian, religion, christians, bible, catholic, christianity, are, islam, commandments, churches | |
13 | violence, insurrection, protests, riots, capitol, january, antifa, was, blm, riot | |
14 | news, media, cnn, msnbc, the, journalism, propaganda, outlets, bias, journalists | |
15 | polls, polling, pollsters, election, betting, trump, is, voters, data, polled | |
19 | vote, voting, party, election, volunteer, candidate, blue, in, will, are | |
democrats_ republicans | 0 | biden, he, and, trump, joe, was, have, if, are, with |
1 | to, is, for, it, as, with, have, woman, just, out | |
2 | inflation, economy, prices, that, tax, for, are, trump, tariffs, all | |
4 | kamala, to, trump, it, harris, as, biden, have, vote, will | |
8 | he, convicted, trump, to, case, trial, judge, jury, prison, court | |
9 | men, trans, are, they, it, gender, black, be, woman, as | |
13 | border, bill, it, immigrants, in, immigration, wall, are, republicans, asylum |
Subreddit Pair | Topic Index | Topic Keywords |
xbox_ playstation | 0 | pro, ps6, upgrade, price, gen, console, years, performance, sony, difference |
1 | reddit, post, what, re, comments, opinion, google, wrong, downvoted, don | |
2 | discs, copies, buy, license, digitally, games, download, store, sales, internet | |
3 | platinum, trophies, achievements, plat, platinums, platinumed, getting, hours, enjoy, games | |
4 | ssd, storage, 2tb, expansion, hdd, 4tb, heatsink, install, seagate, ps5 | |
5 | xbox, gen, console, games, 360, 4k, disc, upgrade, performance, storage | |
6 | ps2, ps1, playstation, nes, owned, sega, memories, n64, atari, snes | |
7 | xbox, exclusives, microsoft, platform, consoles, sony, market, nintendo, platforms, exclusivity | |
8 | gamepass, month, subscription, xbox, games, conversion, buy, gp, price, sales | |
9 | he, wife, kids, dad, together, likes, gift, play, happy, roblox | |
10 | horizontal, overheating, airflow, cooling, vents, ps5, ventilation, exhaust, cabinet, thermal | |
11 | cod, mw3, mw2, warfare, battlefield, warzone, multiplayer, campaigns, modern, bo2 | |
12 | sale, price, friday, full, discount, buy, 60, deals, waiting, games | |
13 | pc, gaming, ps5, consoles, build, pcs, gpu, performance, windows, steam | |
14 | boss, difficulty, fight, level, difficult, enemies, mode, parry, game, hours | |
15 | female, she, characters, protagonist, gender, yasuke, white, japanese, samurai, censored | |
16 | region, psn, vpn, dlc, accounts, steam, sony, uk, requirement, eu | |
17 | router, wifi, ethernet, network, modem, mbps, dns, 5ghz, hotspot, fiber | |
18 | drift, stick, controllers, controller, sticks, dualsense, fix, left, launch, joystick | |
19 | forza, cars, horizon, turismo, motorsport, motorstorm, arcade, motorfest, racer, sim | |
leagueoflegends _DotA2 | 0 | ult, kit, shes, range, winrate, zoe, champion, lane, nerfed, mid |
1 | viktor, lore, characters, show, jayce, zaun, ekko, ionia, episodes, hextech | |
2 | he, worlds, year, player, top, bad, lck, been, performance, mid | |
3 | smurfs, smurfing, accounts, mmr, valve, alt, emerald, ban, matchmaking, level | |
4 | dota, friends, life, fun, games, moba, time, addiction, addicted, much | |
5 | post, read, comments, argument, opinion, downvotes, downvote, discussion, thread, comprehension | |
6 | chat, mute, voice, muting, comms, pinging, toxicity, typing, use, mic | |
10 | he, nerfed, nerfs, kit, lane, champ, mid, ult, buff, patch | |
11 | support, ult, mid, lane, she, adc, lux, jungle, champion, mage | |
12 | fps, gpu, pc, issue, steam, dota, ryzen, reconnect, server, ssd | |
15 | baron, draft, they, pick, dragon, game, geng, call, rumble, drakes | |
17 | reports, 12k, system, toxic, griefing, communication, overwatch, commends, comm, scores | |
18 | mouse, camera, keyboard, hotkeys, buttons, press, cursor, hotkey, screen, bind | |
atheism_ Christianity | 0 | homosexuality, homosexual, sin, lgbtq, lgbt, homosexuals, bible, sinful, heterosexual, leviticus |
1 | post, comment, conversation, here, argument, don, op, read, have, reply | |
2 | abortion, fetus, abortions, woman, choice, medical, unborn, conception, is, care | |
4 | wives, roles, equal, church, misogyny, paul, she, feminism, authority, misogynistic | |
5 | trans, gender, transgender, dysphoria, identity, intersex, medical, transitioning, cis, transgenderism | |
6 | israel, hamas, palestinians, palestine, gaza, genocide, israelis, idf, zionism, conflict | |
7 | islam, muslims, muslim, islamic, sharia, religions, islamophobia, europe, quran, iran | |
8 | atheist, school, raised, religion, believed, catholic, parents, religious, became, wasn | |
9 | slavery, slaves, slave, bible, israelites, exodus, leviticus, servitude, enslaved, testament | |
10 | evolution, species, apes, darwin, macroevolution, ancestor, creationism, organisms, scientists, abiogenesis | |
11 | science, scientific, scientists, scientist, religion, hypothesis, theories, einstein, physics, newton | |
12 | catholics, protestants, catholicism, protestant, pope, orthodoxy, apostolic, protestantism, denominations, bishops | |
13 | music, songs, metal, bands, rap, listening, satanic, genre, hymns, taylor | |
14 | rapture, end, tribulation, raptured, earth, matthew, generation, time, angels, soon | |
15 | hell, torment, place, gehenna, heaven, death, hades, eternity, sheol, soul | |
16 | schools, education, teach, teachers, curriculum, oklahoma, commandments, religions, district, classrooms | |
18 | dress, clothing, hijab, modesty, woman, dressing, covering, nakedness, muslim, nudity | |
19 | religion, religions, people, control, society, world, power, organized, we, masses |
Subreddit Pair | Topic Index | Topic Keywords |
exmuslim_islam | 0 | age, aisha, puberty, child, girl, nine, muhammad, pedophilia, marrying, dolls |
1 | israel, hamas, palestinians, palestine, palestinian, israelis, zionist, idf, zionists, zionism | |
2 | marry, marriage, relationship, he, married, muslim, family, convert, want, date | |
3 | she, muslim, relationship, family, tell, but, will, marry, convert, want | |
4 | hijab, women, cover, niqab, clothing, modesty, naked, muslim, she, hijabi | |
5 | homosexuality, lgbt, lgbtq, queer, homosexual, homophobic, straight, are, gays, sin | |
6 | hindus, indian, bangladesh, caste, hindutva, pakistanis, bangladeshi, bjp, bengali, culture | |
7 | woman, rights, islam, feminism, feminist, muslim, misogynistic, misogyny, feminists, is | |
9 | sin, forgive, forgiveness, repent, allah, repentance, forgiven, forgives, sinning, repenting | |
10 | islam, left, leaving, muslim, started, about, it, believe, convert, didn | |
11 | post, troll, account, comments, sub, argument, debate, reply, read, arguments | |
12 | universe, god, existence, exist, infinite, evidence, dependent, believe, argument, creation | |
13 | fasting, ramadan, eat, water, days, fasts, health, weight, up, if | |
14 | alcohol, drink, drinking, addiction, smoking, haram, alcoholic, weed, nicotine, harmful | |
15 | arabic, language, translations, quran, read, speak, learning, qur, tafsir, arab | |
16 | hadith, hadiths, bukhari, quran, scholars, narration, fabricated, books, authenticity, qur | |
GetMotivated_ getdisciplined | 0 | sleep, bed, wake, alarm, morning, night, waking, hours, sleeping, work |
1 | phone, media, apps, app, screen, scrolling, reddit, instagram, youtube, tiktok | |
2 | porn, addiction, masturbation, is, that, with, sexual, as, have, watching | |
3 | weight, eat, diet, and, healthy, protein, foods, calorie, body, meals | |
4 | gym, workout, exercise, do, it, week, day, feel, working, like | |
5 | adhd, with, medication, diagnosed, can, meds, like, or, get, be | |
6 | thank, for, sharing, it, post, to, and, words, hear, advice | |
7 | read, books, reading, atomic, life, was, habits, this, changed, helped | |
8 | friends, people, social, with, meet, be, talk, group, like, new | |
9 | weed, smoking, smoke, quit, nicotine, quitting, smoked, years, turkey, vape | |
10 | goals, tasks, task, do, list, it, time, day, work, small | |
11 | life, age, young, old, 20s, of, time, are, 40, 30 | |
13 | her, relationship, he, with, will, but, ex, love, for, time | |
14 | quote, is, man, life, he, be, what, will, quotes, from | |
15 | year, was, myself, of, been, 2024, have, years, when, be | |
16 | study, studying, exam, hours, time, if, break, exams, this, minutes | |
17 | post, comment, this, op, advice, it, re, don, like, understand | |
18 | addiction, addictions, drugs, addicted, with, are, can, but, life, will | |
19 | women, dating, date, are, be, relationship, will, that, girl, rejection | |
DecidingToBeBetter _howtonotgiveafuck | 0 | job, at, life, college, degree, it, get, career, be, age |
1 | her, to, that, the, relationship, is, with, in, do, cheating | |
2 | thank, for, much, appreciate, proud, sharing, all, advice, words, keep | |
3 | phone, media, app, apps, use, tiktok, and, facebook, scrolling, screen | |
6 | ugly, looks, attractive, appearance, and, beauty, women, of, yourself, re | |
10 | parents, family, mom, dad, mother, that, he, their, was, be | |
19 | thoughts, meditation, mind, thought, thinking, brain, intrusive, mindfulness, think, be |
Subreddit Pair | Topic Index | Topic Keywords |
realmadrid_Barca | 0 | he, was, ball, injury, player, season, been, goals, as, goal |
1 | xavi, coach, laporta, manager, season, he, club, team, stay, has | |
2 | ref, var, refs, foul, offside, penalty, referee, referees, negreira, madrid | |
3 | we, half, game, score, chances, goal, team, goals, defense, up | |
4 | ballon, award, vini, won, messi, votes, awards, carvajal, deserved, euros | |
5 | mbappe, mbappé, psg, madrid, kylian, ronaldo, player, will, world, and | |
6 | vini, vinicius, neymar, him, player, brazil, jr, has, but, fans | |
7 | mods, post, comments, sub, reddit, twitter, mod, banned, users, discussion | |
9 | kits, logo, shirt, jerseys, nike, authentic, design, crest, replica, badge | |
10 | pedri, gavi, fdj, role, midfield, midfielder, injured, pivot, play, gundo | |
11 | kroos, modric, toni, luka, modrić, retire, midfield, midfielder, midfielders, season | |
12 | contract, sell, clause, salary, million, pay, fee, transfer, loan, value | |
16 | yamal, messi, age, 17, kid, talent, him, young, player, will | |
18 | madrid, fans, barca, sub, real, support, hate, club, comments, barcelona | |
warriors_lakers | 0 | he, shot, shooting, defense, was, season, has, ball, can, and |
1 | klay, thompson, warriors, was, but, bench, contract, in, with, you | |
2 | kerr, minutes, klay, steph, coaching, this, vets, lineups, curry, when | |
3 | ad, lebron, center, he, ball, defense, play, can, big, offense | |
6 | contract, value, trade, million, salary, pay, cap, player, option, paid | |
7 | refs, foul, calls, fouls, officiating, ball, fouled, replay, lakers, bounds | |
11 | draft, son, nba, 55th, lebron, bron, his, james, be, lakers | |
12 | reddit, post, read, comments, re, opinion, what, said, internet, downvotes | |
16 | coach, coaches, coaching, fired, hire, hc, nba, riley, assistants, hiring | |
18 | fans, kobe, laker, lakers, lebron, hate, sub, media, haters, hater | |
apple_Android | 0 | watches, health, garmin, tracking, apple, wrist, smartwatch, apnea, fitbit, and |
1 | ai, features, will, learning, generative, apple, machine, cloud, iphone, it | |
2 | wallet, nfc, payment, tap, digital, contactless, qr, banking, app, debit | |
3 | carplay, tesla, infotainment, automotive, vehicles, android, manufacturers, bmw, toyota, evs | |
5 | airpods, headphones, buds, pair, earbuds, bose, audio, cancellation, sony, beats | |
6 | camera, cameras, processing, quality, photography, sensors, pixel, iphone, telephoto, phones | |
8 | android, iphone, google, phones, apps, os, switch, apple, samsung, back | |
9 | pixel, pixels, samsung, phones, 8a, xl, android, nexus, hardware, camera | |
11 | foldable, folding, foldables, folds, phones, tablet, slab, hinge, unfolded, screens | |
14 | apps, developers, malware, security, android, appstore, alternative, ios, review, apple | |
16 | chrome, browser, safari, firefox, browsers, webkit, mozilla, google, eu, blink | |
17 | imessage, whatsapp, sms, messaging, telegram, messenger, texting, android, texts, iphone | |
18 | comment, opinion, comments, argument, post, downvoted, read, know, thread, troll | |
19 | airtag, network, tags, trackers, pebblebee, bluetooth, google, wallet, locate, ping | |
linux_windows | 0 | linux, windows, use, os, my, it, desktop, users, but, about |
2 | comment, post, this, people, reddit, what, re, read, know, like | |
7 | windows, win11, w11, win10, upgrade, microsoft, it, w10, ui, eol | |
10 | mac, macos, apple, macbook, linux, macs, os, hardware, laptop, air | |
12 | office, libreoffice, excel, onlyoffice, libre, word, microsoft, openoffice, docs, version | |
13 | laptop, laptops, ram, dell, thinkpad, cpu, thinkpads, intel, hardware, ssd | |
15 | keyboard, mouse, ctrl, shortcuts, gestures, layout, touchpad, shortcut, trackpad, button | |
19 | partition, ssd, drives, partitions, ntfs, hdd, ssds, install, sata, gparted |
Subreddit Pair | Topic Index | Topic Keywords |
antiwork_ WorkReform | 0 | rent, housing, homes, property, houses, mortgage, landlords, apartment, income, market |
1 | insurance, healthcare, medicare, universal, companies, deductible, medicaid, hospitals, premiums, claims | |
2 | hours, overtime, hourly, salaried, salary, shifts, 10, per, schedule, workweek | |
3 | taxes, income, taxed, wealth, taxing, billionaires, government, irs, stocks, wealthy | |
4 | ai, automation, robots, art, jobs, technology, automated, automate, machines, artists | |
5 | mcdonald, burger, wendy, franchise, restaurants, prices, fries, chipotle, burgers, menu | |
6 | interviews, hiring, applications, applicants, resumes, process, interviewing, recruiters, interviewer, tests | |
7 | union, unions, dues, unionize, unionized, benefits, contract, unionizing, company, employees | |
8 | billionaires, wealth, billionaire, billion, wealthy, people, millionaires, world, society, hoarding | |
9 | unemployment, fired, employment, employer, termination, fire, severance, department, claim, attorney | |
10 | ceo, ceos, shareholders, profits, shareholder, companies, business, investors, executives, shares | |
11 | boomers, generation, boomer, millennials, generations, older, millennial, parents, generational, genx | |
12 | she, boss, job, was, manager, tell, herself, it, sounds, has | |
13 | argument, read, wrong, what, comments, conversation, arguing, thread, response, opinion | |
14 | kamala, voters, democrats, bernie, election, dems, democratic, pelosi, dnc, voting | |
15 | pto, unlimited, vacation, hours, accrued, company, policy, week, pay, request | |
16 | office, home, hybrid, work, remotely, from, working, company, productive, commute | |
17 | raise, raises, increase, pay, inflation, job, salary, boss, than, months | |
18 | minimum, wage, 25, federal, wages, inflation, increase, hr, prices, workers | |
19 | break, unpaid, minutes, hours, lunches, clock, shift, min, law, work | |
news_ conspiracy | 0 | covid, vaccine, vaccines, flu, vaccinated, pandemic, mrna, disease, vaccination, polio |
1 | russia, ukraine, putin, nato, war, ukrainians, nuclear, europe, invaded, zelensky | |
2 | comment, your, argument, facts, point, conversation, wrong, lol, read, evidence | |
3 | phones, data, google, ads, app, 5g, devices, android, privacy, iphone | |
4 | bible, church, christian, christianity, religious, christians, religions, catholic, believe, sin | |
5 | trans, gender, transgender, dysphoria, children, being, their, lgbtq, cis, is | |
6 | conspiracy, conspiracies, sub, theories, theory, theorists, subreddit, theorist, believe, political | |
7 | gun, guns, firearms, amendment, shootings, laws, militia, firearm, regulated, nra | |
8 | police, cops, officer, officers, training, law, protect, shooting, gun, duty | |
9 | assassination, bullet, attempt, shots, head, snipers, shooting, president, blood, the | |
10 | border, immigration, immigrants, migrants, illegals, immigrant, wall, undocumented, republicans, migrant | |
11 | biden, trump, debate, dementia, president, joe, election, democrats, obama, who | |
12 | abortion, abortions, fetus, roe, woman, rights, embryos, states, law, choice | |
13 | insurance, healthcare, companies, ceo, medicare, aca, claims, medicaid, premiums, private | |
14 | drugs, smoking, nicotine, marijuana, tobacco, addiction, meth, vape, cigarettes, opioids | |
15 | war, terrorism, terrorist, civilians, terrorists, ww3, wars, iraq, bombing, world | |
16 | jury, trial, judge, convicted, prosecution, court, defendant, prosecutors, verdict, convict | |
17 | education, schools, teachers, college, degree, colleges, university, universities, district, admissions | |
18 | climate, co2, fossil, emissions, earth, scientists, atmosphere, environmental, are, science | |
19 | parents, children, parent, bullying, bullies, school, bullied, bully, adults, family | |
Parenting _childfree | 0 | bed, nap, bedtime, baby, crib, naps, hours, months, sleeps, routine |
1 | surgery, vasectomy, procedure, hysterectomy, iud, insurance, sterilization, periods, pill, pregnancy | |
2 | eat, food, eating, foods, meal, meals, eats, fat, cook, protein | |
3 | gap, sibling, siblings, apart, age, each, another, baby, having, sister | |
4 | cat, pet, pets, animal, love, my, have, are, vet, don | |
5 | names, call, nickname, mama, grandma, his, nicknames, change, mr, use | |
6 | party, birthday, parties, cake, birthdays, friends, invites, family, host, attend | |
7 | she, wants, abortion, if, will, it, decision, is, tell, for | |
8 | gifts, gift, christmas, toys, presents, birthday, things, giving, box, items | |
12 | trip, disney, vacation, travel, vacations, beach, fun, hotel, family, park | |
13 | books, read, reading, library, letters, phonics, reader, dyslexia, learning, reads | |
14 | post, op, comments, read, advice, opinion, re, downvoted, troll, thread | |
15 | mil, grandparents, mom, parents, family, relationship, husband, grandparent, grandkids, visit | |
17 | friends, friend, friendship, friendships, new, with, group, kids, hang, she | |
19 | religion, church, religious, christian, catholic, religions, bible, christianity, agnostic, christians |
Subreddit Pair | Topic Index | Topic Keywords |
science_ askphilosophy | 0 | studies, research, scientific, sample, data, article, size, correlation, scientists, published |
1 | philosophy, plato, philosophers, read, philosopher, socrates, books, academic, history, science | |
3 | determinism, deterministic, compatibilism, freedom, responsibility, control, libertarian, choose, universe, causal | |
4 | conservatives, election, democrats, liberal, republican, politics, democratic, conservatism, government, sides | |
5 | racism, race, crime, racial, poverty, whites, cops, races, gangs, immigrants | |
9 | consciousness, brain, physicalism, qualia, idealism, identity, dualism, self, states, body | |
10 | comment, response, point, didn, understand, sorry, appreciate, op, post, this | |
11 | abortion, abortions, fetus, rights, roe, argument, babies, antinatalism, autonomy, procreation | |
science_ philosophy | 3 | capitalism, economy, tax, capitalist, income, socialism, system, profit, government, society |
11 | trans, gender, transgender, dysphoria, identity, cis, puberty, intersex, surgery, transitioning | |
12 | your, comment, re, argument, point, discussion, responding, read, understand, sorry | |
13 | meat, vegan, vegetarian, vegans, veganism, diets, animals, vegetarians, foods, chicken | |
15 | ai, human, intelligence, data, it, machine, robots, robot, agi, learning | |
environment_ climateskeptics | 0 | ev, evs, electric, vehicles, are, that, they, vehicle, charging, battery |
1 | climate, change, science, is, scientists, about, people, what, warming, believe | |
2 | meat, cows, eat, vegan, animal, beef, methane, agriculture, farmers, cattle | |
3 | solar, wind, panels, renewables, electricity, coal, are, turbines, renewable, but | |
4 | florida, rise, insurance, beach, tide, will, change, climate, property, flood | |
5 | she, to, kamala, trump, is, taylor, fracking, vote, if, climate | |
8 | heat, summer, temperatures, degrees, weather, winter, year, days, record, deaths | |
10 | china, emissions, us, capita, world, coal, countries, are, india, ccp | |
11 | oil, fossil, fuels, companies, industry, big, are, gas, exxon, money | |
12 | your, what, comment, troll, don, are, question, have, read, know | |
13 | science, scientific, scientists, the, study, consensus, peer, what, studies, research | |
16 | nuclear, reactors, waste, energy, solar, plants, build, wind, are, more | |
keto_vegan | 1 | potassium, sodium, magnesium, electrolytes, mg, citrate, chloride, ketoade, cramps, supplements |
3 | calories, macros, app, tracking, cronometer, carbs, calculator, per, weigh, body | |
4 | b12, supplements, vitamins, multivitamin, supplementation, yeast, deficiencies, supplementing | |
5 | bring, birthday, family, party, restaurant, dinner, vegan, make, invited, guests | |
7 | ibs, constipation, psyllium, bloating, digestive, stool, probiotics, microbiome, crohn, bacteria | |
12 | cholesterol, ldl, statins, triglycerides, lipid, keto, arteries, elevated, fats, dr | |
15 | thank, appreciate, journey, yourself, sharing, advice, proud, happy, congratulations, helpful | |
carnivore_ vegetarian | 1 | it, meat, they, is, vegetarian, but, just, have, eat, what |
2 | air, pan, cook, fryer, steak, sear, oven, sous, iron, cooking | |
4 | milk, dairy, cream, cheese, yogurt, oat, butter, kefir, coffee, raw | |
Teachers_ homeschool | 4 | covid, immune, flu, pandemic, masks, illness, vaccines, wash, air, system |
5 | math, calculator, memorization, fractions, division, teach, basic, memorize, calculators, tables | |
7 | religion, bible, religious, church, commandments, beliefs, religions, christianity, muslim, atheist | |
8 | reading, read, phonics, dyslexia, letters, spelling, dyslexic, word, program, blending | |
9 | post, comment, comments, op, what, sub, response, opinion, point, read | |
14 | ai, using, writing, write, generated, essay, students, can, prompt, plagiarism | |
16 | parents, children, parenting, parent, people, trauma, adults, think, be, life | |
18 | summer, week, year, august, june, thanksgiving, school, february, month, during |
Subreddit Pair | Topic Index | Topic Keywords |
personalfinance_ wallstreetbets | 0 | wendy, food, me, chicken, chipotle, like, meal, starbucks, pizza, ass |
1 | inflation, cut, recession, economy, cpi, data, gdp, jobs, banks, government | |
2 | tesla, elon, musk, ev, vehicles, robotaxi, teslas, waymo, uber, charging | |
3 | cars, vehicle, maintenance, toyota, financing, buy, honda, dealership, cash, engine | |
4 | comment, post, this, sub, regarded, advice, appreciate, do, like, sorry | |
7 | job, degree, jobs, college, career, hours, salary, field, pay, skills | |
10 | taxes, irs, withholding, refund, 1099, filing, w2, deductions, cpa, married | |
electricvehicles_ regularcarreviews | 2 | comment, post, reddit, people, read, what, re, article, news, wrong |
5 | winter, heating, heater, cabin, battery, ac, weather, temperatures, preconditioning, climate | |
8 | pedal, regen, braking, brakes, accelerator, friction, speed, control, pads, manual | |
10 | trump, biden, democrats, party, election, speech, gop, voters, politics, media | |
16 | truck, trucks, towing, pickups, beds, hauling, cabs, hitch, payload, big | |
19 | cybertruck, ct, truck, cybertrucks, trucks, tesla, r1t, towing, silverado, vehicle | |
simpleliving_ UnethicalLifeProTips | 1 | noise, speakers, sound, headphones, hearing, bluetooth, neighbors, earplugs, noises, sounds |
5 | comment, post, joke, read, downvoted, re, word, sorry, grammar, argument | |
7 | cult, mormon, religion, scientology, churches, mormons, missionaries, jehovah, cults, satanic | |
11 | thank, sharing, appreciate, wish, love, post, beautiful, words, journey, happy | |
15 | clothes, wardrobe, wool, clothing, thrift, brands, outfits, stores, thrifting, items | |
17 | ads, spotify, plex, vpn, hulu, android, subscriptions, torrent, services, piracy | |
classicalmusic_ electronicmusic | 0 | album, it, on, love, with, do, we, just, time, live |
1 | electronic, album, house, techno, and, stuff, from, albums, music, check | |
3 | spotify, cds, streaming, classical, music, artists, it, vinyl, app, radio | |
abortion_prolife | 0 | your, it, to, this, with, weeks, feel, just, will, know |
17 | her, to, it, with, but, can, will, know, support, baby |
Domain | Subreddit Pair | #shared topics | ||||
Subreddit A | Subreddit B | |||||
Gender | AskWomen vs. AskMen | 16 | 384 | 255 | 18 | 18 |
Feminism vs. MensRights | 15 | 360 | 232 | 36 | 36 | |
AskFeminists vs. MensLib | 12 | 288 | 188 | 63 | 63 | |
Politics | AskALiberal vs. AskConservatives | 20 | 480 | 307 | 0 | 0 |
Liberal vs. Conservative | 16 | 384 | 256 | 36 | 36 | |
democrats vs. republicans | 7 | 168 | 108 | 117 | 9 | |
Gaming | xbox vs. playstation | 20 | 480 | 315 | 0 | 0 |
leagueoflegends vs. DotA2 | 13 | 312 | 208 | 63 | 63 | |
Religion | atheism vs. Christianity | 18 | 432 | 267 | 18 | 18 |
exmuslim vs. islam | 16 | 384 | 238 | 18 | 18 | |
Self-improvement | GetMotivated vs. getdisciplined | 19 | 456 | 300 | 9 | 9 |
DecidingToBeBetter vs. howtonotgiveafuck | 7 | 168 | 112 | 99 | 99 | |
Sports | realmadrid vs. Barca | 14 | 336 | 224 | 36 | 36 |
warriors vs. lakers | 10 | 240 | 160 | 45 | 45 | |
Technology | apple vs. Android | 14 | 336 | 219 | 54 | 54 |
linux vs. windows | 8 | 192 | 128 | 99 | 99 | |
social issues | antiwork vs. WorkReform | 20 | 480 | 320 | 0 | 0 |
News | news vs. conspiracy | 20 | 480 | 308 | 0 | 0 |
Parenting | Parenting vs. childfree | 15 | 360 | 233 | 45 | 45 |
Science | science vs. askphilosophy | 8 | 192 | 127 | 108 | 99 |
science vs. philosophy | 5 | 120 | 80 | 90 | 54 | |
Environment | environment vs. climateskeptics | 12 | 288 | 192 | 45 | 45 |
Diet | keto vs. vegan | 7 | 168 | 112 | 90 | 99 |
carnivore vs. vegetarian | 3 | 72 | 46 | 81 | 99 | |
Education | Teachers vs. homeschool | 8 | 192 | 128 | 90 | 63 |
Finance | personalfinance vs. wallstreetbets | 7 | 168 | 110 | 108 | 108 |
Cars | electricvehicles vs. regularcarreviews | 6 | 144 | 96 | 90 | 90 |
lifestyles | simpleliving vs. UnethicalLifeProTips | 6 | 144 | 92 | 90 | 90 |
Music | classicalmusic vs. electronicmusic | 3 | 72 | 48 | 144 | 9 |
Abortion | abortion vs. prolife | 2 | 48 | 32 | 36 | 153 |
Appendix C Human Evaluation
Appendix D Steering via In-context learning
Model | Vanilla | Out-of-topic Few-shot | Subreddit Identifier | In-topic Few-shot |
Qwen2.5-3B-Instruct | 0.493 | 0.487 | 0.52 | 0.575 | 0.586 | ||
Qwen2.5-7B-Instruct | 0.519 | 0.547 | 0.561 | 0.605 | 0.613 | ||
Qwen2.5-14B-Instruct | 0.517 | 0.532 | 0.57 | 0.607 | 0.610 | ||
Qwen2.5-32B-Instruct | 0.532 | 0.573 | 0.587 | 0.639 | 0.645 | ||
Qwen2.5-72B-Instruct | 0.527 | 0.573 | 0.591 | 0.632 | 0.641 | ||
Llama-3.2-3B-Instruct | 0.421 | 0.319 | 0.330 | 0.340 | 0.366 | ||
Llama-3.1-8B-Instruct | 0.485 | 0.455 | 0.475 | 0.432 | 0.555 | ||
Llama-3.3-70B-Instruct | 0.549 | 0.582 | 0.607 | 0.643 | 0.646 | ||
Mistral-7B-Instruct-v0.3 | 0.480 | 0.466 | 0.511 | 0.642 | 0.554 | ||
Claude-3.5-Haiku | 0.529 | 0.323 | 0.563 | 0.511 | 0.516 | ||
Claude-3.7-Sonnet | 0.543 | 0.536 | 0.620 | 0.602 | 0.598 | ||
Deepseek-v3 | 0.316 | 0.261 | 0.314 | 0.306 | 0.311 | ||
gpt-4o-mini | 0.545 | 0.585 | 0.609 | 0.629 | 0.631 |
Model | 12-shot | 24-shot | 36-shot |
Qwen2.5-3B-Instruct | 0.575 | 0.582 | 0.576 |
Qwen2.5-7B-Instruct | 0.605 | 0.617 | 0.613 |
Qwen2.5-32B-Instruct | 0.639 | 0.640 | 0.639 |
Qwen2.5-72B-Instruct | 0.632 | 0.630 | 0.626 |
Llama-3.2-3B-Instruct | 0.340 | 0.330 | 0.377 |
Llama-3.1-8B-Instruct | 0.549 | 0.561 | 0.554 |
Llama-3.3-70B-Instruct | 0.643 | 0.643 | 0.645 |
Deepseek-v3 | 0.306 | 0.321 | 0.308 |
Model name in our paper | Model card in HuggingFace/Anthropic/DeepSeek/OpenAI |
Qwen2.5-3B-Instruct | Qwen/Qwen2.5-3B-Instruct |
Qwen2.5-7B-Instruct | Qwen/Qwen2.5-7B-Instruct |
Qwen2.5-14B-Instruct | Qwen/Qwen2.5-14B-Instruct |
Qwen2.5-32B-Instruct | Qwen/Qwen2.5-32B-Instruct |
Qwen2.5-72B-Instruct | Qwen/Qwen2.5-72B-Instruct |
Llama-3.2-3B-Instruct | meta-llama/Llama-3.2-3B-Instruct |
Llama-3.1-8B-Instruct | meta-llama/Llama-3.1-8B-Instruct |
Llama-3.3-70B-Instruct | meta-llama/Llama-3.3-70B-Instruct |
Mistral-7B-Instruct-v0.3 | mistralai/Mistral-7B-Instruct-v0.3 |
Claude-3.5-Haiku | claude-3-5-haiku-20241022 |
Claude-3.7-Sonnet | claude-3-7-sonnet-20250219 |
Deepseek-v3 | deepseek-chat |
gpt-4o-mini | gpt-4o-mini-2024-07-18 |
Model | Abortion |
Technology | Politics | Gender | Religion | Diet | Music |
Qwen2.5-3B | 0.562 | 0.641 | 0.536 | 0.519 | 0.551 | 0.614 | 0.677 | 0.500 | 0.638 | ||||
Qwen2.5-7B | 0.750 | 0.653 | 0.591 | 0.565 | 0.570 | 0.655 | 0.658 | 0.458 | 0.677 | ||||
Qwen2.5-14B | 0.719 | 0.653 | 0.597 | 0.578 | 0.594 | 0.685 | 0.703 | 0.500 | 0.631 | ||||
Qwen2.5-32B | 0.812 | 0.703 | 0.654 | 0.615 | 0.597 | 0.687 | 0.734 | 0.521 | 0.697 | ||||
Qwen2.5-72B | 0.688 | 0.681 | 0.654 | 0.627 | 0.601 | 0.695 | 0.728 | 0.458 | 0.694 | ||||
Llama-3.2-3B | 0.281 | 0.353 | 0.360 | 0.306 | 0.305 | 0.372 | 0.373 | 0.417 | 0.473 | ||||
Llama-3.1-8B | 0.500 | 0.572 | 0.507 | 0.520 | 0.529 | 0.588 | 0.633 | 0.479 | 0.629 | ||||
Llama-3.3-70B | 0.688 | 0.678 | 0.640 | 0.642 | 0.597 | 0.711 | 0.709 | 0.500 | 0.684 | ||||
Mistral-7B-v0.3 | 0.500 | 0.541 | 0.499 | 0.505 | 0.498 | 0.586 | 0.595 | 0.521 | 0.602 | ||||
Claude-3.5-Haiku | 0.594 | 0.556 | 0.496 | 0.510 | 0.502 | 0.535 | 0.532 | 0.417 | 0.583 | ||||
Claude-3.5-Sonnet | 0.688 | 0.644 | 0.611 | 0.581 | 0.560 | 0.661 | 0.627 | 0.458 | 0.636 | ||||
Deepseek-v3 | 0.312 | 0.375 | 0.303 | 0.286 | 0.308 | 0.319 | 0.310 | 0.229 | 0.311 | ||||
gpt-4o-mini | 0.750 | 0.669 | 0.671 | 0.605 | 0.609 | 0.695 | 0.715 | 0.583 | 0.689 |
Model | Cars | Environment | Gaming | News | Parenting | Finance | Sports | Science | Lifestyles | Education |
Qwen2.5-3B | 0.542 | 0.583 | 0.576 | 0.588 | 0.567 | 0.600 | 0.544 | 0.580 | 0.620 | 0.594 |
Qwen2.5-7B | 0.552 | 0.661 | 0.595 | 0.584 | 0.605 | 0.627 | 0.576 | 0.570 | 0.598 | 0.625 |
Qwen2.5-14B | 0.552 | 0.583 | 0.610 | 0.633 | 0.622 | 0.609 | 0.573 | 0.565 | 0.587 | 0.578 |
Qwen2.5-32B | 0.583 | 0.630 | 0.621 | 0.640 | 0.639 | 0.636 | 0.609 | 0.604 | 0.641 | 0.586 |
Qwen2.5-72B | 0.604 | 0.677 | 0.621 | 0.653 | 0.618 | 0.645 | 0.589 | 0.618 | 0.609 | 0.633 |
Llama-3.2-3B | 0.385 | 0.307 | 0.296 | 0.318 | 0.300 | 0.291 | 0.344 | 0.333 | 0.370 | 0.336 |
Llama-3.1-8B | 0.552 | 0.531 | 0.509 | 0.562 | 0.588 | 0.600 | 0.497 | 0.536 | 0.543 | 0.625 |
Llama-3.3-70B | 0.562 | 0.661 | 0.642 | 0.636 | 0.627 | 0.636 | 0.609 | 0.604 | 0.663 | 0.648 |
Mistral-7B-v0.3 | 0.510 | 0.516 | 0.533 | 0.545 | 0.545 | 0.564 | 0.518 | 0.536 | 0.630 | 0.547 |
Claude-3.5-Haiku | 0.500 | 0.526 | 0.459 | 0.549 | 0.494 | 0.464 | 0.482 | 0.444 | 0.554 | 0.500 |
Claude-3.5-Sonnet | 0.615 | 0.620 | 0.587 | 0.620 | 0.588 | 0.555 | 0.544 | 0.633 | 0.587 | 0.633 |
Deepseek-v3 | 0.229 | 0.286 | 0.319 | 0.308 | 0.339 | 0.236 | 0.286 | 0.295 | 0.304 | 0.281 |
gpt-4o-mini | 0.552 | 0.682 | 0.621 | 0.672 | 0.618 | 0.673 | 0.591 | 0.609 | 0.685 | 0.633 |
Appendix E Steering via Supervised Finetuning
Figure 14 compares steerability under supervised finetuning against in-context learning, focusing on five open-weight models from the Qwen (3B, 7B, and 14B) and Llama (3.2-8B and 3.1-8B) families.
Contrary to expectations, in-context learning consistently outperforms finetuning across most models. For instance, Qwen-2.5-14B achieves 0.607 with in-context learning and only 0.565 after finetuning; Llama-3.1-8B shows a similar drop from 0.555 to 0.509. This suggests that our demonstration set , while sufficient for prompt-based adaptation, may not be large or diverse enough to fully train model weights. However, there are two exceptions. First, Llama-3.2-3B shows improved performance under finetuning (0.377 vs. 0.340), indicating that low-capacity models may benefit more from parameter updates than from relying on in-context learning alone; second, Qwen-2.5-7B narrows the gap between finetuning and prompting, showing similar performance in both settings. This may indicate that mid-size models can partially internalize alignment signals from modest-scale data.
The gap between in-context and finetuned performance may be further explained by data sparsity. While each community C has 200–500 demonstration pairs, these span multiple topics and rhetorical patterns, making it hard for a model to generalize from limited supervision. In contrast, prompting enables models to reason from concrete, topical examples directly.